How much 0.17 HCI is required to make 250 ml of solution with a pH of 1.79.


1
Expert's answer
2022-03-13T15:20:31-0400

[H+] = 10-pH = 10-1.79 = 0.01622 M

n(H+) = n(HCl) = 0.01622 mol/l × 0.250 l = 0.004055 mol

V(solution) = n(HCl)/M(HCl) = 0.004055 mol / 0.17 mol/l = 0.02385 l = 23.85 ml